if a 90.6 ml sample of gas has a partial pressure of 763.75 mmHg at 23.00 xelcius how many moles of gas are present
1
Expert's answer
2016-07-13T05:42:02-0400
PV = nRT, or
n = RT/PV

Than
n = 8.314*(273+23)/90.6*10-6*133.32*763.75 = 266.76 mol